SAT verbal scores are normally distributed with a mean of 489 and a standard deviation of 93. Use the empirical rule (also called 68-95-99.7 rule) to determine what percentage of the scores lie. a) between 210 and 675. b). Above 675?
Answer #1

Solution :

Given that,

= 489

= 93

Using Empirical rule,

(a)

P( - 3< X < + 3) = 99.7%

P(489 - 3 * 93 < X < 489 + 3 * 93) = 99.7%

P(210 < X < 675) = 99.7%

Answer= 99.7%

(b)

P(X > 675) = 1 - P(X < 675) = 0.15%